OpenStudy (anonymous):

Which unit rate is equivalent to 15 miles per gallon?

Mehek (mehek14):

any options?

OpenStudy (anonymous):

yes give me a sec

OpenStudy (anonymous):

Which unit rate is equivalent to 15 miles per gallon? thirty miles over three gallons three gallons over thirty miles four gallons over sixty miles sixty miles over four gallons

Mehek (mehek14):

so which one do you think is equal to \(\dfrac{15}{1}\)

OpenStudy (anonymous):

thats what i was having a hard time with i couldn't decide whether it was a or b

OpenStudy (anonymous):

alright so if you have 15/1=15 which choice do you believe might equal to 15?

OpenStudy (anonymous):

umm a

Mehek (mehek14):

simplify \(\dfrac{30}{3}\) and you get \(\dfrac{10}{1}\)

Mehek (mehek14):

so not equal to \(\dfrac{15}{1}\)

Mehek (mehek14):

so it's not going to be A

OpenStudy (anonymous):

ok

Mehek (mehek14):

and it's not going to be B or C because they flip it

OpenStudy (anonymous):

alright so if you put each of the answers as a fraction and divide you will find your answer

Mehek (mehek14):

so check D \(\dfrac{60}{4}\)

Mehek (mehek14):

if you divide \(\dfrac{60}{4}\) by 4, you get \(\dfrac{15}{1}\)

OpenStudy (anonymous):

so simplify so i tried simplifying every thing else and i got b

Mehek (mehek14):

B is just like A just flipped so for A you got \(\dfrac{10}{1}\), for B it's just flipped = \(\dfrac{1}{10}\)

OpenStudy (anonymous):

thanks

Mehek (mehek14):

so it's not B